How can we determine a compound have which type of hybridization

Grade:9

1 Answers

Arun
5 years ago

First, write down the expanded structure of the compound correctly, with all bonds shown.

Then count the number of atoms connected to the carbon of interest to you.

If the carbon is connected to four other atoms, its hybridisation is sp3 (easy to remember: number of atoms = sum of the superscripts; 1 for s and 3 for p; 1+3 = 4).

If the carbon is connected to three other atoms, its hybridisation is sp2 (easy to remember: number of atoms = sum of the superscripts; 1 for s and 2 for p; 1+2 = 3).

If the carbon is connected to two other atoms, its hybridisation is sp (easy to remember: number of atoms = sum of the superscripts; 1 for s and 1 for p; 1+1 = 2).

Examples: (Unfortunately in the Quora text editor, I cannot draw expanded structures with all bonds shown; so we will use our imagination and just count the atoms attached to the carbon atoms)

Methane → CH4 → four H atoms on the C → hybridisation is sp3.

Ethane → CH3—CH3 → first carbon has fouratoms on it (3 H atoms and 1 C atom) → hybridisation of first carbon is sp3. Same argument for the second carbon also; therefore hybridisation of second carbon is also sp3.

Ethene (or ethylene) → CH2=CH2 → first carbon has three atoms on it (2 H atoms and 1 C atom) → hybridisation of first carbon is sp2. Same argument for the second carbon also; therefore hybridisation of second carbon is also sp2.

Ethyne (or acetylene) → CH=CH (its a triple bond between the C atoms! How do I draw a triple bond in Quora ?) → first carbon has twoatoms on it (1 H atom and 1 C atom) → hybridisation of first carbon is sp. Same argument for the second carbon also; therefore hybridisation of second carbon is also sp.

Propene → CH3-CH=CH2 → first carbon has four atoms on it (3 H atoms and 1 C atom on the right) → hybridisation of first carbon is sp3. Second carbon has three atoms on it (1 H atom and 2 C atoms, left and right) → hybridisation of second carbon is sp2. Third carbon has three atoms on it (2 H atoms and 1 C atom on the left) → hybridisation of third carbon is sp2.